OpenStudy (raiders88):

what is the value of 2x^-2y^-2 for x=3 and y=-2

OpenStudy (raiders88):

plz help

OpenStudy (raiders88):

@welshfella

OpenStudy (welshfella):

plug in the values: = 2 * 3^-2 * 2 * (-2)^-2 negative index = 1 / positive index = 2 2 -- * --- 3^2 (-2)^2

OpenStudy (welshfella):

- can you continue?

OpenStudy (raiders88):

@jess9626

OpenStudy (welshfella):

what is 3^2 ? and (-2)^2 ?

OpenStudy (welshfella):

3^2 = 3* 3 right?

OpenStudy (welshfella):

gotta go

OpenStudy (anonymous):

do you still need help or did welshfella help you understand it.

OpenStudy (raiders88):

i need the answer now

OpenStudy (raiders88):

plz

OpenStudy (raiders88):

@deana99

OpenStudy (raiders88):

@misty1212

OpenStudy (deana99):

I cant help u i'm not smart @raiders88

OpenStudy (raiders88):

thanks anyways :(

OpenStudy (raiders88):

@LynFran

OpenStudy (raiders88):

@amistre64

OpenStudy (raiders88):

plz help

OpenStudy (raiders88):

i just need the answer cause im runnin out of time

OpenStudy (raiders88):

please

OpenStudy (raiders88):

@brittanydosey

OpenStudy (lynfran):

1/18

OpenStudy (raiders88):

thx

OpenStudy (lynfran):

2(3)^-2*(-2)^-2 2(1/9)*(1/4) 2/9*1/4 2/38 1/18